1:3/4 = 2/3:x what is x
Answer by stanbon(75887)

1:3/4 = 2/3:x what is x
--------------
1/(3/4) = (2/3)/x
Cross-multiply to get:
---
x = (3/4)(2/3)
x = 1/2
===============
Cheers,
Stan H.